OpenStudy (anonymous):

A boat's crew rowed 10.5 miles downstream, with the current, in 1.5 hours. The return trip upstream, against the current, covered the same distance, but took 3.5 hours. Find the crew's rowing rate in still water and the rate of the current.

OpenStudy (anonymous):

what is 4 lol the crew's rowing rate in still water or the rate of current?

OpenStudy (anonymous):

oh ok thanks now how do i get the rowing rate in still water?

OpenStudy (anonymous):

yes i think so.

OpenStudy (anonymous):

no it is work on the computer the answer is not given till after.

OpenStudy (anonymous):

Dear i m so sorry. will amend my solution. Downstream = 10.5/1.5 = 7 miles/hr Upstream = 10.5 / 3.5 = 3 miles/hr Difference in speed = 7-3 = 4 miles/hr Now, divide 4 by 2 we get 2 miles/hr speed of current therefore speed of boat = 7 - 2 = 5 miles/hr in still water once again really sorry.

OpenStudy (anonymous):

it is fine lol so still water is 5 and the rate of current is 2? just making sure

OpenStudy (anonymous):

yes it is see (7miles/hr)*(1.5hrs) = 10.5 miles (3miles/hr)*(3.5hrs) = 10.5 miles so current speed 2 miles/ hr speed of boat = 5 miles/hr

OpenStudy (anonymous):

ok thanks